OpenStudy (anonymous):

WILL FAN AND MEDAL Which of the following choices are equivalent to the expression below? choose all that apply. x^3/5 A(x^3)^1/5 B (x^5)^1/3 C (5sqrt(x))^3 D (3sqrt(x))^5 E 5sqrt(x^3) F 3sqrt(x^5)

OpenStudy (anonymous):

@dtuniyants

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

there is a formula. \(\LARGE\color{black}{ \color{blue}{A}^{\color{green}{B}/\color{red}{C}} =(\sqrt[\color{red}{C}]{\color{blue}{A}})^\color{green}{B} }\)

OpenStudy (solomonzelman):

well, it would be same to also say, \(\LARGE\color{black}{ \color{blue}{A}^{\color{green}{B}/\color{red}{C}} =(\sqrt[\color{red}{C}]{\color{blue}{A}^\color{green}{B}}) }\)

OpenStudy (anonymous):

So its a c and d right?

OpenStudy (anonymous):

Sorry I mean E

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

well, it is 5th

OpenStudy (solomonzelman):

5th root.

OpenStudy (anonymous):

So yes it's E

OpenStudy (anonymous):

And A and C

OpenStudy (solomonzelman):

\(\large\color{black}{ \sqrt[5]{x^3} }\) is correct.

OpenStudy (solomonzelman):

I lost connection, that's why I couldn't post it.

OpenStudy (anonymous):

What about a and c

OpenStudy (solomonzelman):

yeah they are also correct.

OpenStudy (solomonzelman):

B D & F are wrong, though...

OpenStudy (anonymous):

Ok thank you